In a non flow process there is heat transfer loss of 1055kJ and internal energy increase of 210kJ determine the work transfer and state whether the process is expansion or compression
1
Expert's answer
2021-10-05T14:33:01-0400

"Q =-1055kJ\\
∆U = 210kJ\\
W =\ ?"


"∆U = Q- W\\
210= -1055-W\\
W = 1265\ kJ"


Since work done is positive, there is an increase in volume.